# Keeping Score: A Quantitative Analysis of How the CHI Community Appreciates Its Milestones

### v2

@@ Abstract.
milestones, CHI, citations, forgetting curves, bibliometric analysis, Matthew effect, quantitative analysis, meta-science, meta-HCI.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713464isbn: 979-8-4007-1394-1/25/04
Figure 1. Forgetting curves (Reisz et al., 2022) for the Proceedings of the ACM Conference on Human Factors in Computing Systems (CHI) from 1981 to 2024. A forgetting curve plots the number of citations CHI papers make to prior papers published during the years plotted on the x-axis. The primary y-axis represents the citations in these publications over time, adjusted for overall growth in the number of publications in the proceedings, and scaled logarithmically for comparison. Each line represents the forgetting curve for papers published in the year following the right-most point of the line. For instance, the curve at the bottom is the forgetting curve for the 2024 CHI proceedings. The peaks in the forgetting curves visually indicate the presence of “milestones”, when read along a vertical line from top to bottom. In other words, several proceedings contain particularly many references to milestones published in that year. For forgotten milestones, this visible peak is lost over time. Here, ‘forgetting’ refers to a decrease in citations over time. We can see how the proceedings of the last few years are distributing their citations more uniformly, with less spikes and dips visible in the curves. This is indicative of the community’s divided attention outside what was in earlier years considered as its key milestones. Independent of the forgetting curves, the secondary y-axis depicts the number of citations from all CHI papers ever published to the top-30 milestones most cited at CHI (see Table 1) per their publication years (x-axis). For instance, Braun and Clarke’s milestone in the year 2006 (Braun and Clarke, 2006) received many citations and the adoption of this milestone in the CHI community is visible as a growing peak in the forgetting curves. The fact that there were no proceedings in 1984 contributes to fluctuations in the older curves.
[…]
Forgetting curves for citations in the Proceedings of the ACM CHI Conference on Human Factors in Computing Systems from 1981 to 2024.
@@ 1. Introduction
Human-Computer Interaction (HCI) is highly interdisciplinary and characterized by constant evolution and changes (Fuchsberger et al., 2021; Bødker, 2015). Recent changes include, to name but a few, an overall growth in the number of scientific papers published each year (Fortunato et al., 2018; Bornmann et al., 2021) and an acceleration of pace in science with mounting career and publication pressures (Edwards and Roy, 2017; Oppenlaender, 2025). In addition, sudden technological advances can shift focus and topics in HCI research. One such technological advance is the recent progress in large auto-regressive language models, such as OpenAI’s ChatGPT (Bubeck et al., 2023; Vaswani et al., 2017). The increasing integration of language models into HCI research reflects broader trends in methodological acceleration and automation (Thomas et al., 2024). While novel AI tools can streamline certain tasks, over-reliance on such tools and their misuse has led to concerns about prompt-hacking (akin to p-hacking) (Morris, 2024) and shallow inquiries that prioritize speed over depth (Thomas et al., 2024; Kobak et al., 2024; Gray, 2024; Liang et al., 2024; Academ-AI, 2024; Arawjo, 2024). As HCI evolves, it is critical to reflect on the types of contributions that drive long-lasting impact, particularly amid shifting research norms and topics. In light of this dynamic context, it is important to explore what we value in HCI and to understand how the CHI community interacts with its intellectual heritage.
The CHI community has a history of citing and valuing its intellectual heritage. For instance, the “Guide to a Successful Paper” at CHI ’16 expressively called authors to “make sure to cite prior work […] in the relevant area” and works “that have had a major influence on your own work” (ACM SIGCHI, 2016). Letting others build on one’s work has been quoted to be “the entire purpose of a CHI Paper” and “lack of references” was identified as “a frequent cause for complaint – and low rating – by reviewers” (ACM SIGCHI, 2016). These sentiments help steer CHI, highlighting the community’s commitment to acknowledging relevant prior work. The emphasis on recognizing intellectual heritage highlights the importance of understanding which works have shaped the field of HCI. However, beyond subjective appreciation, a more systematic approach is needed to quantify the influence of these impactful works.
@@
We present the results of a quantitative analysis of 11,542 articles published in the Proceedings of the ACM CHI Conference on Human Factors in Computing Systems (1981–2024, with the exception of 1984). Our citation curve analysis examines the CHI community’s current attention span and the community’s selective memory. The majority of milestone citations point 5–15 years back in time, suggesting a “sweet spot” of relevance for papers published at CHI. As expected, we also find that the community milestones make diverse types of contributions. Yet, a comparison with existing well-known research contribution taxonomies (Wobbrock and Kientz, 2016; ACM SIGCHI, 2024), reveals that popular CHI contribution types (such as prototypes and systems), but also other contribution types (such as surveys, datasets, and replications), are absent in the CHI community’s milestones. This highlights how some highly popular contribution types—such as prototypes and systems—all play a role but seldom leave a long-lasting historical mark. We present the 30 milestones most cited by the CHI community in this paper, and we make a dataset of the 300 most cited milestones available in the spirit of Open Science. [^1] Link to Zenodo will be provided after anonymous peer review. We further present a case study on CHI’s “super milestone”, Braun and Clarke’s Thematic Analysis (Braun and Clarke, 2006), that continues to enjoy extraordinary popularity at CHI. Finally, we present the Milestone Coefficient as a metric for assessing research impact of milestones. This metric offers a more nuanced understanding of what constitutes a milestone beyond mere citation count. We discuss our findings in light of related literature and provide our reflections on how the community remembers, forgets, and curates past knowledge.
Our research contributes to meta research on HCI literature (“meta-HCI”) (Oppenlaender et al., 2025), particularly focusing on the CHI community, as a not fully representative but important community in the field of HCI.
@@ 2. Related Work
While the HCI field is buzzing with activity, as demonstrated by the well-visited and growing CHI Conference, surprisingly little research and self-reflection on HCI research is being published at CHI. For example, the HCI field’s top conference currently has no subcommittee for meta-scientific investigations. This contrasts with the field’s critical and self-reflective tradition (Sengers et al., 2005; Schön, 1992). Such self-reflective and meta-scientific contributions are critically important to advance the HCI field as a whole, and in the following we present prior meta-research on HCI and CHI.
@@ 2.4. Forgetting Curves
A citation curve plots the number of citations to past papers from papers published in a given year. For instance, if a paper published in the year 2024 cites a paper published in 1950, this counts as one citation to the 1950 paper. Citation curves provide a visual indication of the attention span of authors. On the one hand, a flat citation curve indicates a long community memory where many old papers are being cited in a given proceedings year. On the other hand, a citation curve with a steep and right-leaning peak indicates a short community memory where old papers are quickly being forgotten. However, one factor that comes into play is the number of articles published each year. This number has been increasing in recent years, thereby introducing a bias to the insights that we can derive from citation curves.
Forgetting curves (Reisz et al., 2022) address this problem. A forgetting curve is a growth-adjusted citation curve. For all references in a given proceedings year, the forgetting curve looks at what prior years have been cited and counts citations to these years (i.e., the citation curve). The forgetting curve is calculated as follows:
(1) $\bar{c}_{t_{i}t_{j}}=\frac{N_{t_{j}}}{N}\sum_{t_{j}}c_{t_{i}t_{j}}$
where $N_{t_{j}}$ is the number of papers published in year $t_{j}$, $N$ is the total number of papers published in all years, and $c_{t_{i}t_{j}}$ is the citation curve (Reisz et al., 2022). The sum $\sum_{t_{j}}c_{t_{i}t_{j}}$ is the total number of citations from papers published in year $i$ to papers published in year $j$. To mitigate bias due to growth in publications over the years, the forgetting curve is weighted with the relative number of publications. This weighting factor in formula (1) together with the citation counts can be interpreted as an expectancy value. Valleys in the forgetting curve will form when citations to a past year under-perform compared to the expected value, and peaks will indicate above average performance. See the work by Reisz et al. (Reisz et al., 2022) for a more detailed explanation of citation curves and forgetting curves.
@@ 3.2.1. Milestone definition and identification
As mentioned in Section 2.2, there is no standard method for identifying milestone papers and different approaches have been used in prior work. In our work, the papers with the highest total number of citations from CHI papers (over all CHI proceedings) are considered the milestones of the CHI community. More specifically, we adopt the approach by Reisz et al. (Reisz et al., 2022) and define the top 30 most cited publications as milestones in our initial analysis. As in Reisz et al., we acknowledge this choice is arbitrary. However, to solidify our analysis we later extend our analysis to the top-100, top-300, and top-3000 papers cited in the CHI community, finding that the results are consistent for larger groups of milestones. We further define “super milestones” as high performing milestones which have received more than 0.1% (1 in 1000) of all citations in at least one proceedings year. We derived this threshold empirically from visual inspection of citations over time (cf. Figure 2). While this threshold is also arbitrary, it serves as a meaningful decision boundary for discussing different types of milestones in our analysis.
Note that in our analysis and throughout this whole article, milestones are not limited to being published at CHI, but can also appear in other publication venues. A milestone can be a conference paper, journal article, or whole book. For simplicity, we refer to them as ‘milestones’ or ‘milestone papers’ in the remainder of this work. Further note that the citation counts presented in our analysis (cf. Table 1) differ significantly from those that include all citations to a milestone, such as citation counts reported by Google Scholar. Our focus in this work rests exclusively on the CHI community and how it cites its milestones.
To identify the milestones, we parse the reference texts and extract the publication year and the title with an iteratively developed set of over 100 regular expressions and logical checks. We apply basic normalization to the paper’s title, including lowercasing as well as removal of spaces, line breaks, and special characters. By concatenating the year and normalized title into a unique ID, we can robustly identify papers by this ID (though, see Section 5.5 for limitations of this approach). To answer RQ1, we can then analyse the citations from CHI papers to the uniquely identified papers.
@@ 3.2.2. Milestone types and contributions
To identify the milestone contributions, we inductively coded the key contributions of the top-30 milestones most cited in the CHI community. To this end, the first author read the milestones’ title and abstract and iteratively coded the milestones. The codes were discussed in two meetings with the co-authors to reach consensus. This approach ensured consistency and alignment in the interpretation of the data. One code was assigned to each milestone paper. We acknowledge that a milestone may make more than one contribution. However, our coding aimed to capture the key contribution of the community’s top-performing milestones, answering the question of “what is the milestone most known and cited for?” The identified set of contributions was then mapped to existing contribution types by Wobbrock and Kientz (Wobbrock and Kientz, 2016) and ACM SIGCHI (ACM SIGCHI, 2024).
@@ 3.3. Forgetting Curve Plots
We plot forgetting curves for the available CHI Proceedings years. The construction of forgetting curves is straight-forward from the scraped data, using the formula given in Section 2.4. The forgetting curves include citations to papers published at CHI and other venues. Therefore, the curves reach far back in time, but for the purpose of analyzing CHI, we limit the plot to our time frame of analysis (1981–2024). On the secondary axis (in Figure 1), we plot the total number of citations to the 30 most cited milestones in their respective publication year as vertical bars. The bars help in visually identifying some of the milestone peaks in the forgetting curves.
@@ 3.4. Milestone Coefficient
We define the milestone coefficient as a measure for the growing or diminishing impact of a paper relative to other papers published within a research community. The aim of this metric is to classify papers on a continuous open-ended “milestone spectrum.” We design the metric to amplify differences between milestones that were popular once, but are being forgotten, and milestones that continue to grow in citations. Our aim with this metric is to provide a tool for analyzing scientific papers cited by a research community. The metric goes beyond a discrete typology of milestones and instead uses a continuous spectrum to measure the success of a paper in a research community. To this end, we first define the trend of the citation series $c$ of a paper:
@@
where $x$ are the years in the citation series $c$ and $\alpha$ is an exponential factor that adjusts for the fact that citation accrual follows a power law. For estimating the parameter $\alpha$, we analyze all citations ever made in the CHI proceedings. For each cited paper, we construct the citation series (citations over the years) and fit a powerlaw distribution to these series. This results in an estimate of $\alpha=3.63$. TCI is the first component of the Milestone Coefficient. The second component is the Normalized Citation Impact (NCI) which provides a measure of a paper’s citation performance relative to all other citations to papers in a proceedings years $x$. NCI is defined as the sum of this relative impact:
(2) $\text{NCI}_{j}=\sum_{x}\frac{\text{citations}_{jx}}{\sum_{i}\text{citations}_{ix}}$
@@
$\text{MC}=\text{SIGN}~{}\times~{}\sqrt{\text{TCI}^{2}~{}\times~{}\text{NCI}^{2}}$

### v3

@@ Abstract.
milestones, CHI, citations, forgetting curves, Matthew effect, quantitative analysis, bibliometrics, meta-science, meta-HCI.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713464isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing Empirical studies in HCIccs: General and reference Surveys and overviews
Figure 1. Forgetting curves (Reisz et al., 2022) for the Proceedings of the ACM Conference on Human Factors in Computing Systems (CHI) from 1981 to 2024. A forgetting curve plots the number of citations from CHI papers to prior papers published during the years plotted on the x-axis. The primary y-axis represents the citations in these publications over time, adjusted for overall growth in the number of publications in the proceedings, and scaled logarithmically for comparison. Each line represents the forgetting curve for papers published in the year following the right-most point of the line. For instance, the curve at the bottom is the forgetting curve for the 2024 CHI proceedings. The peaks in the forgetting curves visually indicate the presence of “milestones”, when read along a vertical line from top to bottom. In other words, several proceedings contain particularly many references to milestones published in that year. For forgotten milestones, this visible peak is lost over time. Here, ‘forgetting’ refers to a decrease in citations over time. We can see how the proceedings of the last few years are distributing their citations more uniformly, with less spikes and dips visible in the curves. This is indicative of the community’s divided attention outside what was in earlier years considered as its key milestones. Independent of the forgetting curves, the secondary y-axis depicts the number of citations from all CHI papers ever published to the top-30 milestones most cited at CHI (see Table 1) per their publication years (x-axis). For instance, Braun and Clarke’s milestone in the year 2006 (Braun and Clarke, 2006) received many citations and the adoption of this milestone in the CHI community is visible as a growing peak in the forgetting curves. The fact that there were no proceedings in 1984 contributes to fluctuations in the older curves.
@@ 1. Introduction
Human-Computer Interaction (HCI) is highly interdisciplinary and characterized by constant evolution and changes (Fuchsberger et al., 2021; Bødker, 2015). Recent changes include, to name but a few, an overall growth in the number of scientific papers published each year (Fortunato et al., 2018; Bornmann et al., 2021) and an acceleration of pace in science with mounting career and publication pressures (Edwards and Roy, 2017; Oppenlaender, 2025). In addition, sudden technological advances can shift focus and topics in HCI research. One such technological advance is the recent progress in large auto-regressive language models, such as OpenAI’s ChatGPT (Bubeck et al., 2023; Vaswani et al., 2017). The increasing integration of language models into HCI research reflects broader trends in methodological acceleration and automation (Thomas et al., 2024; Pang et al., 2025). While novel AI tools can streamline certain tasks, over-reliance on such tools and their misuse has led to concerns about prompt-hacking (akin to p-hacking) (Morris, 2024) and shallow inquiries that prioritize speed over depth (Thomas et al., 2024; Kobak et al., 2024; Gray, 2024; Liang et al., 2024; Academ-AI, 2024; Arawjo, 2024). As HCI evolves, it is critical to reflect on the types of contributions that drive long-lasting impact, particularly amid shifting research norms and topics. In light of this dynamic context, it is important to explore what we value in HCI and to understand how the CHI community interacts with its intellectual heritage.
The CHI community has a history of citing and valuing its intellectual heritage (Oppenlaender, 2025). For instance, the “Guide to a Successful Paper” at CHI ’16 expressively called authors to “make sure to cite prior work […] in the relevant area” and works “that have had a major influence on your own work” (ACM SIGCHI, 2016). Letting others build on one’s work has been quoted to be “the entire purpose of a CHI Paper” and “lack of references” was identified as “a frequent cause for complaint – and low rating – by reviewers” (ACM SIGCHI, 2016). These sentiments help steer CHI, highlighting the community’s commitment to acknowledging relevant prior work. The emphasis on recognizing intellectual heritage highlights the importance of understanding which works have shaped the field of HCI. However, beyond subjective appreciation, a more systematic approach is needed to quantify the influence of these impactful works.
@@
We present the results of a quantitative analysis of 11,542 articles published in the Proceedings of the ACM CHI Conference on Human Factors in Computing Systems (1981–2024, with the exception of 1984). Our citation curve analysis examines the CHI community’s current attention span and the community’s selective memory. The majority of milestone citations point 5–15 years back in time, suggesting a “sweet spot” of relevance for papers published at CHI. As expected, we also find that the community milestones make diverse types of contributions. Yet, a comparison with existing well-known research contribution taxonomies (Wobbrock and Kientz, 2016; ACM SIGCHI, 2024), reveals that popular CHI contribution types (such as prototypes and systems), but also other contribution types (such as surveys, datasets, and replications), are absent in the CHI community’s milestones. This highlights how some highly popular contribution types—such as prototypes and systems—all play a role but seldom leave a long-lasting historical mark. We present the 30 milestones most cited by the CHI community in this paper. In the spirit of Open Science, we share a dataset of the 300 most cited milestones in the Supplemental Material of this article. We further present a case study on CHI’s “super milestone”, Braun and Clarke’s Thematic Analysis (Braun and Clarke, 2006), that continues to enjoy extraordinary popularity at CHI. Finally, we present the Milestone Coefficient as a metric for assessing research impact of milestones. This metric offers a more nuanced understanding of what constitutes a milestone beyond mere citation count. We discuss our findings in light of related literature and provide our reflections on how the community remembers, forgets, and curates past knowledge.
Our research contributes to meta-research on HCI literature (“meta-HCI”) (Oppenlaender et al., 2025), particularly focusing on the CHI community, which, while not fully representative, is an important community in the field of Human-Computer Interaction.
@@ 2. Related Work
While the field of HCI is buzzing with activity, as evident in the well-visited and growing annual CHI Conference, surprisingly little research and self-reflection on HCI research is being published at CHI. For example, the HCI field’s top conference currently has no subcommittee for meta-scientific investigations. This contrasts with the field’s critical and self-reflective tradition (Sengers et al., 2005; Schön, 1992). Such self-reflective and meta-scientific contributions are critically important to advance the HCI field as a whole, and in the following we present prior meta-research on HCI and CHI.
@@ 2.4. Forgetting Curves
A citation curve plots the number of citations to past papers from papers published in a given year. For instance, if a paper published in the year 2024 cites a paper published in 1950, this counts as one citation to the year 1950. Citation curves provide a visual indication of the attention span of authors. On the one hand, a flat citation curve indicates a long community memory where many old papers are being cited in a given proceedings year. On the other hand, a citation curve with a steep and right-leaning peak indicates a short community memory where old papers are quickly being forgotten. However, one factor that comes into play is the number of articles published each year. This number has been increasing in recent years, thereby introducing a bias to the insights that we can derive from citation curves.
Forgetting curves (Reisz et al., 2022) address this problem. A forgetting curve is a growth-adjusted citation curve. For all references in a given proceedings year, the forgetting curve looks at what prior years have been cited and counts citations to these years (i.e., the citation curve), adjusted for growth in publications. The forgetting curve is calculated as follows:
$\bar{c}_{t_{i}t_{j}}=\frac{N_{t_{j}}}{N}\sum_{t_{j}}c_{t_{i}t_{j}}$
where $N_{t_{j}}$ is the number of papers published in year $t_{j}$, $N$ is the total number of papers published in all years, and $c_{t_{i}t_{j}}$ is the citation curve (Reisz et al., 2022). The sum $\sum_{t_{j}}c_{t_{i}t_{j}}$ is the total number of citations from papers published in year $i$ to papers published in year $j$. To mitigate bias due to growth in publications over the years, the forgetting curve is weighted with the relative number of publications. This weighting factor together with the citation counts can be interpreted as an expectancy value. Valleys in the forgetting curve will form when citations to a past year under-perform compared to the expected value, and peaks will indicate above average performance. See the work by Reisz et al. (Reisz et al., 2022) for a more detailed explanation of citation curves and forgetting curves.
@@ 3.2.1. Milestone definition and identification
As mentioned in Section 2.2, there is no standard method for identifying milestone papers and different approaches have been used in prior work. In our work, the papers with the highest total number of citations from CHI papers (over all CHI proceedings) are considered the milestones of the CHI community. More specifically, we adopt the approach by Reisz et al. (Reisz et al., 2022) and define the top 30 most cited publications as milestones in our initial analysis. As in Reisz et al., we acknowledge this choice is arbitrary. However, to solidify our analysis we later extend our analysis to the top-100, top-300, and top-3000 papers cited in the CHI community, finding that the results are consistent for larger groups of milestones.
We further define “super milestones” as high performing milestones which have received more than 0.1% (1 in 1000) of all citations in at least one proceedings year. We derived this threshold empirically from visual inspection of citations over time (cf. Figure 2). While this threshold is also arbitrary, it serves as a meaningful decision boundary for discussing different types of milestones in our analysis.
Note that in our analysis and throughout this whole article, milestones are not limited to being published at CHI, but can also appear in other publication venues. A milestone can be a conference paper, journal article, or whole book. For simplicity, we refer to them as ‘milestones’ or ‘milestone papers’ in the remainder of this work. Further note that the citation counts presented in our analysis (cf. Table 1) differ significantly from those that include all citations to a milestone, such as citation counts reported by Google Scholar. Our focus in this work rests exclusively on the CHI community and how it cites its milestones (published at CHI or any other venues).
To identify the milestones, we parse each reference text to extract the publication year and the title with an iteratively developed set of over 100 regular expressions and logical checks. We apply basic normalization to the paper’s title, including lowercasing as well as removal of spaces, line breaks, and special characters. By concatenating the year and normalized title into a unique ID, we can robustly identify papers by this ID (though, see Section 5.5 for limitations of this approach). To answer RQ1, we can then analyze the citations from CHI papers to the uniquely identified papers.
@@ 3.2.2. Milestone types and contributions
To identify the milestone contributions, we inductively coded the key contributions of the top-30 milestones most cited in the CHI community. To this end, the first author read the milestones’ title and abstract and iteratively coded the milestones. The codes were discussed in two meetings with the co-author to reach consensus. This approach ensured consistency and alignment in the interpretation of the data. One code was assigned to each milestone paper. We acknowledge that a milestone may make more than one contribution. However, our coding aimed to capture the key contribution of the community’s top-performing milestones, answering the question of “what is the milestone most known and cited for?” The identified set of contributions was then mapped to existing contribution types by Wobbrock and Kientz (Wobbrock and Kientz, 2016) and ACM SIGCHI (ACM SIGCHI, 2024).
@@ 3.3. Forgetting Curve Plots
We plot forgetting curves for the available CHI Proceedings years. The construction of forgetting curves is straight-forward from the collected data, using the formula given in Section 2.4. The forgetting curves include citations to papers published at CHI and other venues. Therefore, the curves reach far back in time, but for the purpose of analyzing CHI, we limit the plot to our time frame of analysis (1981–2024). On the secondary axis (in Figure 1), we plot the total number of citations to the 30 most cited milestones in their respective publication year as vertical bars. The bars help in visually identifying some of the milestone peaks in the forgetting curves.
@@ 3.4. Milestone Coefficient
We define the milestone coefficient as a measure for the growing or diminishing impact of a paper relative to other papers published within a research community. The aim of this metric is to classify papers on a continuous open-ended “milestone spectrum.” We design the metric to amplify differences between milestones that were popular once, but are being forgotten, and milestones that continue to grow in citations. Our aim with this metric is to provide a tool for analyzing highly-cited scientific papers in a research community. The metric goes beyond a discrete typology of milestones and instead uses a continuous spectrum to measure the success of a paper in a research community. To this end, we first define the trend of the citation series $c$ of a paper:
@@
where $x$ are the years in the citation series $c$ and $\alpha$ is an exponential factor that adjusts for the fact that citation accrual follows a power law. For estimating the parameter $\alpha$, we analyze all citations ever made in the CHI proceedings. For each cited paper, we construct the citation series (citations over the years) and fit a powerlaw distribution to these series. This results in an estimate of $\alpha=3.63$. TCI is the first component of the Milestone Coefficient.
The second component is the Normalized Citation Impact (NCI) which provides a measure of a paper’s citation performance relative to all other citations to papers in a proceedings years $x$. NCI is defined as the sum of this relative impact:
$\text{NCI}_{j}=\sum_{x}\frac{\text{citations}_{jx}}{\sum_{i}\text{citations}_{ix}}$
@@
$\text{MC}=\text{SIGN}~{}\cdot~{}\sqrt{\text{TCI}^{2}~{}\cdot~{}\text{NCI}^{2}}$
Figure 2. Relative number of citations to CHI’s top-100 most cited milestones over time. A clustering analysis reveals three types of milestones in CHI: one “super milestone” continues to grow in popularity (red), some past super milestones are fading from the CHI community’s memory (light blue), and some contemporary milestones are receiving continued attention, although at lower levels (blue).